Suppose that a Monopolists market demand is given by P a - Q, a>2, and that the a) Calculate the profit maximising monopoly price and quantity. b) Calculate the price and quantity that arise under perfect competition. [8 marks ] c) Calculate and compare Consumer and Producer Surplus both under monopoly [6 marks ] and perfect competition: what is the Deadweight loss due to Monopoly? Provide a graphical description of the two cases. [16 marks]
Answer #1

Given, P = a - Q

TR = aQ - Q2

MR = a - 2Q

MC = 2

A. Monopolist will maximize profit by satisfying the following relationship

MR = MC

a - 2Q = 2

large 2Q = a - 2

large Q = rac{a - 2}{2}( where a>2)

P = a - Q

large P = a -rac{a - 2}{2}

large P = rac{2a - a + 2}{2}

a P=

Profit maximizing price = (a+2)/2

Profit maximizing quantity = (a-2)/2

B. In case of perfect competition the firm will maximize its profit by equating price with marginal cost

P = MC

a - Q = 2

Profit maximizing quantity, QP = a - 2

Profit maximizing price, PP = a - (a -2) = $ 2 per unit

C. Monopoly

CS A+ R-2 MC ar- 2

Consumer Surplus

large CS = rac{1}{2}*(a-rac{a+2}{2})* rac{a-2}{2}

large CS = rac{(a-2)^2}{8}

Producer surplus

large PS = (rac{a+2}{2}-2)*rac{a-2}{2}

large PS = rac{(a-2)^2}{4}

Dead weight loss

a +2 2 k (a- 2

DIVL = (4-2)2

Perfect Competition

2 LC 0

Consumer Surplus

large CS =rac{1}{2}*(a-2)(a-2)

cs a-2)

Please contact if having any query thank you.